OpenStudy (anonymous):

Solve the equation 1296 x = 6

OpenStudy (anonymous):

x=6/1296 now solve

OpenStudy (saifoo.khan):

easy

OpenStudy (anonymous):

solving x as an Exponential?

OpenStudy (saifoo.khan):

no!

OpenStudy (saifoo.khan):

x = 6/1296

OpenStudy (saifoo.khan):

x = 1/216